OpenStudy (anonymous):

Determine a formula for estimating the volume and floor area of the building. Estimate the effect of the irregular shape of the Tower. The tour guide tells you that the building is 110 stories high and sits on one city block (1/16 mile square). The typical height between floors in a commercial builing is about 12 feet or about 4 meters.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

since when is a square irregular? Area of a square A= S^2 , in meter.. so convert miles to meter. ( MKS standard units) 1/ 16 mile X (1609 meters/1mile ) =100.56m, now A=S^2= 100.56^2= 10112m^2 Volume = A X H Area we have, need height Height= 110 stories X 4 meters = 440 m Volume = A X H = 10112 X 440= 44449280 m^3

OpenStudy (anonymous):

sorry this site doesnt email me notifications for some reason but its the sears tower. its not a perfect shape. its built form is buddled tube or something